A and B are real non-zero 3 × 3 matrices and satisfy the equation (AB)^T + B^(-1) A = 0. Prove that if B is orthogonal then A is antisymmetric.


1
Expert's answer
2021-03-31T16:35:28-0400

Using property of orthogonal matrix:


"B^{-1}=B^T"

we have


"(AB)^T+B^{-1}A=B^TA^T+B^TA"

"=B^T(A^T+A)=0"

Since "B" is non-zero, product is zero when:


"A^T+A=0"

"A=-A^T"

Therefore in this case "A" is antisymmetric.